A chemist prepares a solution of barium acetate
by measuring out
of barium acetate into a
volumetric flask and filling the flask to the mark with water.
Calculate the concentration in
of the chemist's barium acetate solution. Round your answer to
significant digits.
1
Expert's answer
2019-04-01T07:49:50-0400

Solution.

Since it does not say how much barium acetate was taken, we assume that 100 grams of barium acetate were taken and the volumetric flask had a volume of 300 milliliters.

M((CH3COO)2Ba) = 255.43 g/mole

V(solution) = 300 mL

C = m/(M*V)

C = 100/(255.43*0.3) = 1.30 M

Answer:

C = 100/(255.43*0.3) = 1.30 M